OpenStudy (candyme):

The dimensions of the old stage at the concert hall were 30ft wide and 15ft deep. The new stage has a total area of 1000 square ft. The dimensions of the new stage were created by adding the same distance x to the width and the depth of the old stage dimensions. What is the value of x?

OpenStudy (anonymous):

Length * Width = Area The dimensions of the old stage at the concert hall were 30ft wide and 15ft deep => Length = 15 => Width = 30 The dimensions of the new stage were created by adding the same distance x to the width and the depth of the old stage dimensions. => New Length = 15+X => New Width = 30+X The new stage has a total area of 1000 square ft => New Area = 1000 (New Length)*(New Width) = New Area Solve for X.